OpenStudy (anonymous):

two friends can do a piece of work for $960 . one alone can do it in 12 days and other in 16 days. If with assistance of an expert they finish in 4 days what should be the remuneration the expert should get? A)$ 700 B)$ 550 C)$ 400 D)$ 350

OpenStudy (anonymous):

@LastDayWork

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@ganeshie8

OpenStudy (lastdaywork):

Can you find the number of days the expert will take to complete the work on his own ??

OpenStudy (anonymous):

1/12 + 1/16 + 1/x = 1/4 x pert = 9.6 hours by himself or $100 per day. 4 days = $400 for the expert.